A stock has an expected return of 9 percent, its beta is 0.85, and the risk-free rate is 3.6 percent. What must the expected return on the market be?

a)9.95%

b)10.35%

c)6.35%

d)9.46%

e)10.45%
Answer #1

expected return=risk-free rate +Beta*(market rate- risk-free rate )

9=3.6+0.85*(Market rate-3.6)

(9-3.6)=0.85*(Market rate-3.6)

Market rate=(9-3.6)/0.85+3.6

=9.95%(Approx).